# Antiferromagnetism and Tightly Bound Cooper Pairs Induced by Kinetic Frustration

### v1

@@ Appendix D Singlet formation enhancement under hole doping
A minimal four-site model captures the energy reduction in kinetic energy when forming spin singlet state. The model set up on a four-site plaquette:
$\mathcal{H}=-t_{1}\sum_{\langle jl\rangle,\sigma}c_{j\sigma}^{\dagger}c_{l\sigma}-t_{2}\sum_{\langle\langle jl\rangle\rangle,\sigma}c_{j\sigma}^{\dagger}c_{l\sigma}$ (21)
The ground state for different spin sectors are shown in Fig. S9. While at $t_{2}/t_{1}=0$ fully polarized state have the lowest energy when dopping one holes, the lowest spin sectors consistently have the lowest energy for $t_{2}/t_{1}>0.25$ for both single- and two-hole doping, confirming that singlet formation is energetically favored. As expected, the stability of these low-spin states increases with $t_{2}/t_{1}$, supporting the interpretation that KF promotes singlet formation.
A similar trend appears in the extended system. As shown in Fig. S10, increasing $t_{2}/t_{1}$ from 00 to $0.6$ suppresses the strong triplet correlations, initially characterized by $M_{i,j,(k)}\approx 0.37$, and instead stabilizes antiferromagnetic correlations with $M_{i,j,(k)}\approx 0.61$. This AFM correlation persists even in the two-hole case when the holes are doped but remain spatially separated. While there is a modest enhancement of singlet correlations around each hole, it remains limited compared to the contribution arising from spin interactions alone.
In contrast, when two holes are doped adjacently, singlet tendencies are significantly amplified—far exceeding the enhancement expected from spin-spin interactions alone (see Fig.5). Fig. S11 illustrates this cooperative effect, showing an increase in singlet correlations from 0.58 to approximately 0.82, with excellent agreement between ED and DMRG results. The ED data match the DMRG findings closely, with a maximum deviation of only 3% in the singlet correlation measured around the doped holes.
@@
Appendix E Kinetic picture from effective correlated hopping
In this section, we demonstrate that singlet formation around fixed hole positions contributes to kinetic energy enhancement when two holes are close. In the main text, we map the original model’s kinetic energy contribution—which incorporates spin fluctuations in the full Hilbert space—into a two-particle model with hopping that depends on the other particle’s position, known as correlated hopping. With the effective wavefunction $\psi(j,k)$ defined as the square root of the hole-hole correlation function $P_{j,k}$, this reformulation give exactly the same kinetic energy as original full $t-J$ model.
The effective correlated hopping $|\tilde{t}_{jk}(l)|$ will never exceed $|t_{j,k}|$, owing to the Cauchy-Schwarz inequality:
$\langle\Psi_{G}|\bar{n}_{l}\bar{n}_{j}|\Psi_{G}\rangle\langle\Psi_{G}|\bar{n}_{l}\bar{n}_{k}|\Psi_{G}\rangle$ (22)
$=$ $\langle\Psi_{G}|\bar{n}_{l}\sum_{j,\sigma}c_{j\sigma}c_{j\sigma}^{\dagger}|\Psi_{G}\rangle\langle\Psi_{G}|\bar{n}_{l}\sum_{j,\sigma}c_{k\sigma}c_{k\sigma}^{\dagger}|\Psi_{G}\rangle$
$\geq$ $|\langle\Psi_{G}|\bar{n}_{l}c_{k\sigma}c_{j\sigma}^{\dagger}\bar{n}_{l}|\Psi_{G}\rangle|^{2}$
A clear enhancement in both nearest-neighbor and next-nearest-neighbor hopping can be observed around each hole. This is demonstrated using both Hilbert space truncated ED on an $8\times 8$ periodic boundary condition (PBC) cluster and DMRG on an $8\times 16$ cylinder, as shown in Fig. S12.
Figure S12: DMRG-ED comparison for the correlated hopping strength $\tilde{t}_{jk}(l)/t_{jk}$. DMRG calculations use the same parameters as in Fig. 6 in the main text, and ED calculations are performed on an $8\times 8$ cluster with $\lambda=\Delta=1,U=10t_{1}$. (a, c) show vertical/horizontal hopping and (b, d) show diagonal hopping. (a, b) are DMRG results and (c, d) are ED results.
The enhancement in hole correlation patterns directly accounts for the formation of tightly bound Cooper pairs. We demonstrate this using the an equivalent two-particle model, extended to the thermodynamic limit. When two holes approach each other, the hopping amplitude is replaced by values obtained from Hilbert space truncated ED of the Hilbert space with truncation, as shown in Fig. S12 (c, d). When the holes are far apart, we adopt an averaged hopping value among bonds with largest separation allowed from the hole with Hilbert space truncated ED. Values are obtained separately for NN hopping and NNN hopping on different sublattices. The resulting hole-hole correlations of the two-particle model are shown in Fig. S13 (a). The coherence length $\xi$ is extracted by fitting the exponential decay of the Cooper pair wavefunction at large hole separation, using the form:
$|\psi(i,j)|^{2}\sim\frac{1}{r}e^{-\frac{r}{\xi}}$ (23)
where $r$ is the distance between sites $i$ and $j$. As shown in Fig. S13 (b), Correlations are well captured by the fitting with coherence length $\xi=9.6$
@@ Appendix E Kinetic picture from effective correlated hopping
Figure S13: (a) Hole-hole correlation $P_{i,j}=|\psi(i,j)|^{2}$, calculated using correlated hopping amplitudes extracted from ED of the truncated Hilbert space on an $8\times 8$ system, without including the energy differences induced by Ising spin interactions. The calculation is performed with $\lambda=\Delta=1$ and $U=10t_{1}$. Only the central region is visualized, though the full calculation spans a much larger region to approximate the thermodynamic limit. (b) Visualization of $|r_{i}-r_{j}|P_{i,j}$ as a function of $r=|r_{i}-r_{j}|$, based on the data in (a). (c) Hole-hole correlation $P_{i,j}$ calculated with the same hopping enhancement as in (a), but including an attractive interaction of strength $J_{1}/2$ from the Ising spin interaction. (d) Visualization of $|r_{i}-r_{j}|P_{i,j}$ as a function of $r=|r_{i}-r_{j}|$, based on the data in (c).
@@ Appendix F DMRG calculation details
We select a longer length $L_{x}=16$ comparing to $L_{y}$ in order to minimize open boundary effects on hole density distributions. Open boundaries suppress hole density at system edges, reducing the likelihood of hole-hole correlations involving boundary-located holes. This boundary effect accounts for the observed breaking of $C_{4}$ symmetry in correlation patterns.
@@
Finally, we benchmark our DMRG results against exact diagonalization at a larger interaction strength $U=12.0t_{1}$. As shown in Fig. S15, both methods exhibit consistent correlation behavior with no qualitative differences, confirming the robustness of our results at stronger coupling.
@@
Figure S15: ED-DMRG agreement persists at higher interaction strength, validating our computational approach. We compare hole-hole correlations versus distance at $(\lambda,\Delta)=(1,1)$ and $U=12.0t_{1}$. Error bars in DMRG data represent correlation spreads at identical distances due to boundary effects that reduce hole density away from cylinder centers. The excellent agreement confirms robust pairing physics across interaction regimes.

### v2

@@ Appendix D Singlet formation enhancement under hole doping
A minimal four-site model captures the energy reduction in kinetic energy when forming spin singlet state. The ground states for different spin sectors are shown in Fig. S9. While the fully polarized state has the lowest energy in the single-hole sector at $t_{2}/t_{1}=0$, the lowest-spin sectors consistently become energetically favored for $t_{2}/t_{1}>0.25$ in both the one- and two-hole cases, confirming that singlet formation is preferred. As expected, the stability of these low-spin states increases with $t_{2}/t_{1}$, supporting the interpretation that kinetic frustration promotes singlet formation near the holes.
A similar trend is observed in larger lattices. As shown in Fig. S10, increasing $t_{2}/t_{1}$ from 00 to $0.6$ suppresses the strong triplet correlations, initially characterized by $M_{i,j,(k)}\approx 0.37$, and instead stabilizes antiferromagnetic correlations with $M_{i,j,(k)}\approx 0.61$. This AFM correlation persists even in the two-hole case when the holes are doped but remain spatially separated. While there is a modest enhancement of singlet correlations around each hole, it remains limited compared to the contribution arising from spin interactions alone.
In contrast, when two holes occupy NN sites, singlet tendencies are significantly amplified—far exceeding the enhancement expected from spin-spin interactions alone (see Fig.5). Fig. S11 illustrates this cooperative effect, showing an increase in singlet character from 0.58 to approximately 0.82, with excellent agreement between ED and DMRG results. The ED data match the DMRG findings closely, with a maximum deviation of only 3% in the singlet correlation measured around the doped holes.
@@
Appendix E Effective two-body hopping
In this section, we demonstrate that singlet formation around fixed hole positions enhances the kinetic energy when two holes are nearby. In the main text, we map the kinetic energy contribution of the original model—which incorporates spin fluctuations within the full Hilbert space—onto a two-particle model with a two-body hopping term, i.e., a hopping operator that depends explicitly on the position of the other particle. Using the effective wavefunction $\psi(j,k)$, defined as the square root of the hole-hole correlation function $P_{j,k}$, this reformulation accurately reproduces the kinetic energy of the original full $t$-$J$ model.
The effective two-body hopping $|\tilde{t}_{jk}(l)|$ will never exceed $|t_{j,k}|$, owing to the Cauchy-Schwarz inequality:
$\langle\Psi_{G}|\bar{n}_{l}\bar{n}_{j}|\Psi_{G}\rangle\langle\Psi_{G}|\bar{n}_{l}\bar{n}_{k}|\Psi_{G}\rangle\geq|\langle\Psi_{G}|\bar{n}_{l}\sum_{\sigma}c_{k\sigma}c_{j\sigma}^{\dagger}\bar{n}_{l}|\Psi_{G}\rangle|^{2}.$
A clear enhancement of both NN and NNN hopping amplitudes is observed in the vicinity of each hole. This effect is demonstrated using Hilbert space–truncated exact diagonalization on an $8\times 8$ cluster with PBC, as well as DMRG calculations on an $8\times 16$ cylindrical geometry, as shown in Fig. S12.
Figure S12: DMRG-ED comparison for the two body hopping strength $\tilde{t}_{jk}(l)/t_{jk}$. DMRG calculations use the same parameters as in Fig. 6 in the main text, and ED calculations are performed on an $8\times 8$ cluster with $\lambda=\Delta=1,U=10t_{1}$. (a, c) show vertical/horizontal hopping and (b, d) show diagonal hopping. (a, b) are DMRG results and (c, d) are ED results.
The enhancement of the effective two-body hopping amplitude at short hole-hole distances directly accounts for the formation of tightly bound Cooper pairs. We demonstrate this using the effective two-particle model extended to the thermodynamic limit, in which the two-body hopping amplitudes are extracted from our Hilbert space–truncated ED calculations, as shown in Fig. S12(c, d). For large hole separations, we assign an average hopping value based on the bonds for the largest allowed hole-hole distance, as determined from the truncated ED data. Separate values are used for nearest-neighbor and next-nearest-neighbor hopping on different sublattices. The resulting hole-hole correlations from the two-particle model are shown in Fig. S13(a). The coherence length $\xi$ is extracted by fitting the exponential decay of the Cooper pair wavefunction at large hole separation using the form:
$|\psi(i,j)|^{2}\propto\frac{1}{\rho}e^{-\frac{\rho}{\xi}}$ (23)
where $\rho$ is the distance between sites $i$ and $j$. As shown in Fig. S13 (b), Correlations are well captured by the fitting with coherence length $\xi=9.6$
@@ Appendix E Effective two-body hopping
Figure S13: (a) Hole-hole correlation $P_{i,j}=|\psi(i,j)|^{2}$, calculated using two-body hopping amplitudes extracted from ED of the truncated Hilbert space on an $8\times 8$ system, without including the energy differences induced by Ising spin interactions. The calculation is performed with $\lambda=\Delta=1$ and $U=10t_{1}$. Only the central region is shown, though the full calculation spans a much larger region to approximate the thermodynamic limit. (b) Visualization of $|r_{i}-r_{j}|P_{i,j}$ as a function of $\rho=|r_{i}-r_{j}|$, based on the data in (a). (c) Hole-hole correlation $P_{i,j}$ calculated with the same hopping enhancement as in (a), but including an attractive interaction of strength $J_{1}/2$ from the Ising spin interaction. (d) Visualization of $|r_{i}-r_{j}|P_{i,j}$ as a function of ${\rho}$, based on the data shown in panel (c).
@@ Appendix F DMRG calculation details
We select a longer length $L_{x}=16$ relative to $L_{y}$ in order to minimize open boundary effects on hole density distribution. Open boundaries suppress hole density on the edges, reducing the likelihood of hole-hole correlations involving boundary-located holes. This boundary effect accounts for the observed breaking of $C_{4}$ symmetry in correlation patterns.
@@
Finally, we benchmark our DMRG results against exact diagonalization for larger interaction strength $U=12.0t_{1}$. As shown in Fig. S15, both methods exhibit consistent correlation behavior with no qualitative differences, confirming the robustness of our results as a function of increasing $U/t_{1}$.
@@
Figure S15: The agreement between ED and DMRG persists for higher interaction strength $U/t_{1}$, validating our computational approach. We compare the hole density-density correlation function versus distance for $(\lambda,\Delta)=(1,1)$ and $U=12.0t_{1}$. Error bars in the DMRG data represent correlation spreads at identical distances due to boundary effects that reduce hole density away from cylinder centers.